# Multi-Layer Cycle Benchmarking for high-accuracy error characterization

### v1

@@ Appendix D Random noise models
$f^{B}_{IXIII}$ $f^{B}_{IXIXI}f^{B}_{ZXIXZ}$ $+0.50$ $\begin{pmatrix}-0.25\\ +0.25\\ -0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{ZXIII}f^{B}_{ZXZII}\\ f^{B}_{IXIII}f^{B}_{IXZII}\\ f^{B}_{IIIXI}f^{B}_{IIIXZ}\\ \end{pmatrix}$
@@
$f^{G}_{IIIXI}$ $f^{G}_{IXZXZ}f^{G}_{ZXZXI}$ $-0.50$ $\begin{pmatrix}-0.50\\ +0.75\\ +0.25\\ +0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IIZII}f^{G}_{IIZII}\\ f^{G}_{IIZXI}f^{G}_{IIIXI}\\ f^{G}_{IIIXZ}f^{G}_{IIZXZ}\\ f^{G}_{IXZII}f^{G}_{ZXZII}\\ \end{pmatrix}$
@@
$f^{B}_{IXIII}$ $f^{B}_{IXIXI}f^{B}_{ZXIXZ}$ $+0.50$ $\begin{pmatrix}+0.25\\ -0.25\\ -0.25\\ +0.25\\ -0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IXZII}f^{B}_{IXIII}\\ f^{B}_{ZIIXI}f^{B}_{ZIIXZ}\\ f^{B}_{ZXIII}f^{B}_{ZXZII}\\ f^{B}_{ZIIII}f^{B}_{ZIIII}\\ f^{B}_{IIIXZ}f^{B}_{IIIXI}\\ \end{pmatrix}$
@@
$f^{G}_{IIIXI}$ $f^{G}_{IXZXZ}f^{G}_{ZXZXI}$ $-0.50$ $\begin{pmatrix}+0.25\\ +0.25\\ -0.50\\ +0.50\\ -0.25\\ +0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IIIXZ}f^{G}_{IIZXZ}\\ f^{G}_{ZIIXI}f^{G}_{ZIZXI}\\ f^{G}_{IIZII}f^{G}_{IIZII}\\ f^{G}_{IXZII}f^{G}_{ZXZII}\\ f^{G}_{ZIIII}f^{G}_{ZIIII}\\ f^{G}_{IIIXI}f^{G}_{IIZXI}\\ \end{pmatrix}$
@@
$f^{B}_{IXIII}$ $f^{B}_{IXIXI}f^{B}_{ZXIXZ}$ $+0.50$ $\begin{pmatrix}+0.50\\ -0.50\\ -0.25\\ -0.25\\ +0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IXZII}f^{B}_{IXIII}\\ f^{B}_{IIIXI}f^{B}_{IIIXZ}\\ f^{B}_{IXIIZ}f^{B}_{IXZIZ}\\ f^{B}_{ZXIII}f^{B}_{ZXZII}\\ f^{B}_{IIIIZ}f^{B}_{IIIIZ}\\ \end{pmatrix}$
@@
$f^{G}_{IIIXI}$ $f^{G}_{IXZXZ}f^{G}_{ZXZXI}$ $-0.50$ $\begin{pmatrix}+0.25\\ +0.50\\ -0.25\\ +0.25\\ +0.75\\ -0.25\\ -0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IXIIZ}f^{G}_{ZXIIZ}\\ f^{G}_{IXZII}f^{G}_{ZXZII}\\ f^{G}_{ZXIII}f^{G}_{IXIII}\\ f^{G}_{IIIXZ}f^{G}_{IIZXZ}\\ f^{G}_{IIIXI}f^{G}_{IIZXI}\\ f^{G}_{IIIIZ}f^{G}_{IIIIZ}\\ f^{G}_{IIZII}f^{G}_{IIZII}\\ \end{pmatrix}$
@@
$f^{B}_{IXII}$ $f^{B}_{IXXX}f^{B}_{IYYX}$ $+0.50$ $\begin{pmatrix}+0.25\\ -0.50\\ -0.25\\ -0.25\\ -0.25\\ -0.25\\ -0.25\\ +0.50\\ +0.25\\ +0.25\\ +0.50\\ -0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IIXY}f^{B}_{IIXY}\\ f^{B}_{IIXY}f^{B}_{IIYX}\\ f^{B}_{IXIX}f^{B}_{ZXZX}\\ f^{B}_{ZIII}f^{B}_{ZIII}\\ f^{B}_{IIXX}f^{B}_{IIXX}\\ f^{B}_{IYIX}f^{B}_{ZYZX}\\ f^{B}_{IIZI}f^{B}_{IIZI}\\ f^{B}_{IIZX}f^{B}_{IIIX}\\ f^{B}_{ZIZI}f^{B}_{ZIZI}\\ f^{B}_{IXII}f^{B}_{ZXII}\\ f^{B}_{IXII}f^{B}_{ZYII}\\ f^{B}_{ZYII}f^{B}_{IYII}\\ \end{pmatrix}$
@@
$f^{G}_{IIXI}$ $f^{G}_{ZXYY}f^{G}_{ZYXY}$ $-0.50$ $\begin{pmatrix}-0.25\\ +0.25\\ +0.25\\ -0.25\\ +0.25\\ +0.25\\ +0.50\\ -0.25\\ -0.25\\ +0.25\\ +0.50\\ -0.25\\ +0.25\\ -0.50\\ -0.25\\ +0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{ZZII}f^{G}_{ZZII}\\ f^{G}_{IXIY}f^{G}_{IXIY}\\ f^{G}_{IIXY}f^{G}_{ZZXY}\\ f^{G}_{IIYI}f^{G}_{ZIYI}\\ f^{G}_{ZXII}f^{G}_{ZXIZ}\\ f^{G}_{ZYII}f^{G}_{ZYIZ}\\ f^{G}_{IXIX}f^{G}_{IYIY}\\ f^{G}_{IXIZ}f^{G}_{IXII}\\ f^{G}_{ZIII}f^{G}_{ZIII}\\ f^{G}_{IZII}f^{G}_{IZII}\\ f^{G}_{IIXI}f^{G}_{ZIYI}\\ f^{G}_{IXIX}f^{G}_{IXIX}\\ f^{G}_{ZIXI}f^{G}_{IIXI}\\ f^{G}_{IIIY}f^{G}_{IZIY}\\ f^{G}_{IYII}f^{G}_{IYIZ}\\ f^{G}_{IIYY}f^{G}_{ZZYY}\\ \end{pmatrix}$

### v2

@@ Appendix F Random noise models
$f^{B}_{IXIII}$ $f^{B}_{IXIXI}f^{B}_{ZXIXZ}$ +0.50+0.50 $\begin{pmatrix}-0.25\\ +0.25\\ -0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{ZXIII}f^{B}_{ZXZII}\\ f^{B}_{IXIII}f^{B}_{IXZII}\\ f^{B}_{IIIXI}f^{B}_{IIIXZ}\\ \end{pmatrix}$
@@
$f^{G}_{IIIXI}$ $f^{G}_{IXZXZ}f^{G}_{ZXZXI}$ 0.50-0.50 $\begin{pmatrix}-0.50\\ +0.75\\ +0.25\\ +0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IIZII}f^{G}_{IIZII}\\ f^{G}_{IIZXI}f^{G}_{IIIXI}\\ f^{G}_{IIIXZ}f^{G}_{IIZXZ}\\ f^{G}_{IXZII}f^{G}_{ZXZII}\\ \end{pmatrix}$
@@
$f^{B}_{IXIII}$ $f^{B}_{IXIXI}f^{B}_{ZXIXZ}$ +0.50+0.50 $\begin{pmatrix}+0.25\\ -0.25\\ -0.25\\ +0.25\\ -0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IXZII}f^{B}_{IXIII}\\ f^{B}_{ZIIXI}f^{B}_{ZIIXZ}\\ f^{B}_{ZXIII}f^{B}_{ZXZII}\\ f^{B}_{ZIIII}f^{B}_{ZIIII}\\ f^{B}_{IIIXZ}f^{B}_{IIIXI}\\ \end{pmatrix}$
@@
$f^{G}_{IIIXI}$ $f^{G}_{IXZXZ}f^{G}_{ZXZXI}$ 0.50-0.50 $\begin{pmatrix}+0.25\\ +0.25\\ -0.50\\ +0.50\\ -0.25\\ +0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IIIXZ}f^{G}_{IIZXZ}\\ f^{G}_{ZIIXI}f^{G}_{ZIZXI}\\ f^{G}_{IIZII}f^{G}_{IIZII}\\ f^{G}_{IXZII}f^{G}_{ZXZII}\\ f^{G}_{ZIIII}f^{G}_{ZIIII}\\ f^{G}_{IIIXI}f^{G}_{IIZXI}\\ \end{pmatrix}$
@@
$f^{B}_{IXIII}$ $f^{B}_{IXIXI}f^{B}_{ZXIXZ}$ +0.50+0.50 $\begin{pmatrix}+0.50\\ -0.50\\ -0.25\\ -0.25\\ +0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IXZII}f^{B}_{IXIII}\\ f^{B}_{IIIXI}f^{B}_{IIIXZ}\\ f^{B}_{IXIIZ}f^{B}_{IXZIZ}\\ f^{B}_{ZXIII}f^{B}_{ZXZII}\\ f^{B}_{IIIIZ}f^{B}_{IIIIZ}\\ \end{pmatrix}$
@@
$f^{G}_{IIIXI}$ $f^{G}_{IXZXZ}f^{G}_{ZXZXI}$ 0.50-0.50 $\begin{pmatrix}+0.25\\ +0.50\\ -0.25\\ +0.25\\ +0.75\\ -0.25\\ -0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IXIIZ}f^{G}_{ZXIIZ}\\ f^{G}_{IXZII}f^{G}_{ZXZII}\\ f^{G}_{ZXIII}f^{G}_{IXIII}\\ f^{G}_{IIIXZ}f^{G}_{IIZXZ}\\ f^{G}_{IIIXI}f^{G}_{IIZXI}\\ f^{G}_{IIIIZ}f^{G}_{IIIIZ}\\ f^{G}_{IIZII}f^{G}_{IIZII}\\ \end{pmatrix}$
@@
$f^{B}_{IXII}$ $f^{B}_{IXXX}f^{B}_{IYYX}$ +0.50+0.50 $\begin{pmatrix}+0.25\\ -0.50\\ -0.25\\ -0.25\\ -0.25\\ -0.25\\ -0.25\\ +0.50\\ +0.25\\ +0.25\\ +0.50\\ -0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IIXY}f^{B}_{IIXY}\\ f^{B}_{IIXY}f^{B}_{IIYX}\\ f^{B}_{IXIX}f^{B}_{ZXZX}\\ f^{B}_{ZIII}f^{B}_{ZIII}\\ f^{B}_{IIXX}f^{B}_{IIXX}\\ f^{B}_{IYIX}f^{B}_{ZYZX}\\ f^{B}_{IIZI}f^{B}_{IIZI}\\ f^{B}_{IIZX}f^{B}_{IIIX}\\ f^{B}_{ZIZI}f^{B}_{ZIZI}\\ f^{B}_{IXII}f^{B}_{ZXII}\\ f^{B}_{IXII}f^{B}_{ZYII}\\ f^{B}_{ZYII}f^{B}_{IYII}\\ \end{pmatrix}$
@@
$f^{G}_{IIXI}$ $f^{G}_{ZXYY}f^{G}_{ZYXY}$ 0.50-0.50 $\begin{pmatrix}-0.25\\ +0.25\\ +0.25\\ -0.25\\ +0.25\\ +0.25\\ +0.50\\ -0.25\\ -0.25\\ +0.25\\ +0.50\\ -0.25\\ +0.25\\ -0.50\\ -0.25\\ +0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{ZZII}f^{G}_{ZZII}\\ f^{G}_{IXIY}f^{G}_{IXIY}\\ f^{G}_{IIXY}f^{G}_{ZZXY}\\ f^{G}_{IIYI}f^{G}_{ZIYI}\\ f^{G}_{ZXII}f^{G}_{ZXIZ}\\ f^{G}_{ZYII}f^{G}_{ZYIZ}\\ f^{G}_{IXIX}f^{G}_{IYIY}\\ f^{G}_{IXIZ}f^{G}_{IXII}\\ f^{G}_{ZIII}f^{G}_{ZIII}\\ f^{G}_{IZII}f^{G}_{IZII}\\ f^{G}_{IIXI}f^{G}_{ZIYI}\\ f^{G}_{IXIX}f^{G}_{IXIX}\\ f^{G}_{ZIXI}f^{G}_{IIXI}\\ f^{G}_{IIIY}f^{G}_{IZIY}\\ f^{G}_{IYII}f^{G}_{IYIZ}\\ f^{G}_{IIYY}f^{G}_{ZZYY}\\ \end{pmatrix}$